# Implementation and Performance Evaluation of TCP over QUIC Tunnels

### v1

@@ 6.3 Latency Impact Analysis
Sender Receiver Sender Receiver Sender Receiver Sender Receiver
@@
Observations:
  • Throughput Decrease with Latency: Both protocols exhibit the expected inverse relationship between latency and throughput. As latency increases, throughput decreases for both QUIC Tunnel and Plain TCP. While QUIC Tunnel maintains a higher throughput compared to Plain TCP at lower latencies (e.g., 0 ms and 50 ms), the performance gap narrows as latency increases. At higher latencies, both protocols experience significant throughput degradation.
  • Relative Stability of QUIC Tunnel: QUIC Tunnel demonstrates relatively more stable throughput at higher latencies. For instance, at 1000 ms, the throughput of QUIC Tunnel is still notably higher (6.60 Mbps on the sender side, 2.76 Mbps on the receiver side) compared to Plain TCP (0.73 Mbps and 0.61 Mbps, respectively). However, QUIC Tunnel’s performance is not immune to the impact of increasing latency.
  • CPU Usage Variability: CPU usage for both QUIC Tunnel and Plain TCP shows significant variability, likely due to factors such as system load and network conditions. While QUIC Tunnel tends to exhibit somewhat higher CPU usage at certain points (e.g., at 0 ms latency), it does not consistently outperform Plain TCP in terms of CPU usage. The observed differences in CPU usage between the two protocols are subject to random fluctuations and should not be used as a definitive measure of efficiency.
  • Impact of Latency on Plain TCP: Plain TCP shows a more pronounced decrease in throughput as latency increases, with a particularly sharp decline at latencies of 500 ms and beyond. In contrast, QUIC Tunnel’s throughput degradation is less severe under similar conditions, indicating that QUIC’s protocol design offers a degree of resilience to latency.
@@ 6.4 Packet Reordering Tolerance
The out-of-order rate is formally defined as:
[…]
where $N_{\text{out-of-order}}$ represents the number of packets received out of order, and $N_{\text{total}}$ is the total number of packets received. This metric provides insight into the system’s tolerance and recovery from packet reordering during transmission.
The out-of-order delivery tests from Table 4 reveal:
@@
Sender Receiver Sender Receiver Sender Receiver Sender Receiver
@@
And we can find that:
  • QUIC’s advantage in reordering: QUIC can handle packet reordering more effectively due to its stream-based, out-of-order tolerant design.
  • TCP over QUIC doesn’t fully inherit QUIC’s advantages: Despite being encapsulated in QUIC, TCP still relies on its in-order delivery and retransmission mechanisms, which can be affected by reordering.
  • Slightly improved stability over native TCP: TCP over QUIC may show slightly better stability compared to native TCP because QUIC can buffer and reorder packets more efficiently at the transport layer.
@@ 7 Conclusion & Future Work
Further research could explore more forms of QUIC-based network tunneling, such as TCP over WebTransport [22], to further develop the potential of QUIC in the transformation of the network protocol. In addition, TCP over QUIC tunnel can optimize the processing of different application layer protocols and improve the performance of traditional application layer protocols under this tunnel.

### v2

@@ 6.3 Latency Impact
Sndr Rcvr Sndr Rcvr Sndr Rcvr Sndr Rcvr
@@
From Table 3 and Fig. 3:
  • Throughput decreases with latency for both protocols, but the tunnel maintains higher throughput at moderate latencies.
  • At 1000 ms, the tunnel is still clearly ahead (6.60 vs. 0.73 Mbps on sender side).
  • CPU usage varies, with the tunnel often slightly higher, but not consistently better or worse.
@@ 6.4 Packet Reordering Tolerance
Out-of-order rate is defined as:
$\text{Out-of-Order Rate}=\frac{N_{\text{out-of-order}}}{N_{\text{total}}}\times 100\%.$
@@
Sndr Rcvr Sndr Rcvr Sndr Rcvr Sndr Rcvr
@@
As shown in Table 4 and Fig. 4:
  • QUIC tolerates reordering well due to its stream-based design.
  • TCP over QUIC inherits some benefit but remains constrained by TCP’s in-order semantics.
  • The tunnel shows slightly better stability than native TCP under reordering.
Finally, note that both inner TCP and outer QUIC implement congestion control independently. Their interaction may cause inefficiencies such as oscillation, which require further investigation beyond our VM-based tests.
@@ 7 Conclusion & Future Work
Our current evaluation is limited to a single QUIC implementation (Quinn) and VM-based setups. Broader validation across diverse QUIC stacks and real-world WAN environments will further strengthen the conclusions.
Further research could explore more forms of QUIC-based network tunneling, such as TCP over WebTransport [7], to further develop the potential of QUIC in the transformation of the network protocol. In addition, TCP over QUIC tunnel can optimize the processing of different application layer protocols and improve the performance of traditional application layer protocols under this tunnel. Future work should also investigate the congestion control interactions between the inner TCP and the outer QUIC layers, as this remains a fundamental challenge for tunneling designs.
8 Data Availability
The data sets generated and analyzed during the current study are not persistently stored, as all experimental data were dynamically generated during runtime. Due to variations in network conditions, system load, and inherent randomness, storing raw data such as packet captures or instantaneous performance metrics was deemed unnecessary. To ensure reproducibility, all testing methods and the full source code used in this study are publicly available at the GitHub repository: https://github.com/ElaBosak233/quic-tun. Researchers can reproduce the results by building and executing the code under similar or customized network conditions.